OpenStudy (anonymous):

The surface areas of two similar solids are 334 yd2 and 1085 yd2. The volume of the larger solid is 1755 yd3. What is the volume of the smaller solid? A) 300 yd3 B) 974 yd3 C) 810 yd3 D) 1,755 yd3 Thanks!

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

Thanks satellite!

OpenStudy (anonymous):

your ratio of the surface areas is \[\frac{334}{1775}\] that is area and comes with the square. the volume comes with the cube, so your ratio is \[(\sqrt{\frac{334}{1775}})^3=.171\] rounded. multiply \[.171\times 1775=300\] rounded

OpenStudy (anonymous):

awesome, thank you